A horizontal force of 0•80N is Required to pull A 5kg Block Across a table top at a constant speed. With the block initially at rest again. Determine the speed v of the bullet where the bullet is assumed to be embedded in the block.

1
Expert's answer
2022-04-02T10:11:37-0400

"Fd=\frac{mv^2}2,"

"v=\sqrt{\frac{2Fd}m}=\sqrt{0.32d}."